### v1

# Ab initio study of Proximity-Induced Superconductivity in PbTe/Pb heterostructures

Abstract
Semiconductor–superconductor hybrid devices have been proposed as promising platforms for detecting and analyzing Majorana zero modes, which find applications in topological quantum computing. In this work, we solve the Kohn-Sham Density Functional Theory and Bogoliubov–de Gennes equations to describe the normal and superconducting properties of a PbTe/Pb heterostructure. We resolve a proximity–induced superconducting gap on the PbTe side. The hybridization between PbTe and Pb causes the emergence of a soft Bardeen–Cooper–Schrieffer–like superconducting gap. We compute the anomalous charge density in real space, estimating its decay length and showing that the pairing potential is anisotropic, which is a necessary condition for unconventional superconductivity. Contrary to the models that predict Majorana zero modes in these interfaces, we find a significantly large Schottky barrier in the normal state preventing the emergence of zero modes. Our findings strengthen the understanding of the physics governing PbTe/Pb hybrid devices and their viability for Majorana zero modes applications.
keywords:
superconductivity, semiconductor-superconductor heterostructure, BdG equations
1 Introduction
PbTe/Pb nanowires are a promising material platform to host Majorana Zero Modes (MZMs) due to their resilience to charged impurities, high electron mobility, and excellent crystalline quality  1, 2. The intrinsic properties of PbTe (narrow bandgap and large Landé g-factor) combined with the superconducting gap ($\Delta\sim$ 3meV) and relatively high critical temperature ($T_{C}\sim$ 7K) of Pb, make these materials advantageous compared to conventional InAs-Al and InSb-Al nanowires 3, 4.
To host MZMs, a semiconductor–superconductor (SM–SC) interface must meet several key requirements 5, 6: (i) strong spin-orbit coupling and significant Zeeman energy gap ($V_{Z}$) in the semiconductor, (ii) $s$–wave superconductor, (iii) proximity induced superconductivity, and (iv) the heterostructure chemical potential must lie within the SM’s Zeeman gap and the SC gap. When the applied Zeeman field exceeds the critical value $V_{C}=\sqrt{\Delta^{2}+\mu^{2}}$ the system transitions into a topological phase, hosting MZMs at its ends 7, 8, 9.
Cole et al. 10 developed a model to describe the density of states (DOS) in a generic nanowire and investigated how the SC–coupling strength and the height of the interface barrier affect the shape of the SC gap. They identified a hard U–shaped superconducting gap as indicative of a strong SC–coupling regime, where the superconductor overshadows the semiconductor’s electronic properties. Conversely, a soft Bardeen–Cooper–Schrieffer–like gap corresponds to a weak SC–coupling regime, where low-energy states of the HS are predominantly of semiconductor character. Weak SC–coupling is, therefore, desirable for achieving topological superconductivity 11, 12.
In this work, we investigate from first–prinicples proximity effects in a PbTe/Pb heterostructure (HS), focusing on the conditions required for MZMs to emerge. All our findings are obtained using the recently implemented SIESTA-BdG method 13, 14, which solves the Kohn-Sham Density Functional Theory and Bogoliubov–de Gennes equations to describe on the same footing the normal and superconducting properties. We discuss how the electrostatic environment at the interface is critical for the emergence of MZMs. We find significant band bending at the interface, leading to charge transfer from PbTe to Pb, and affecting the position of the chemical potential. In the superconducting state we observe a soft BCS–like SC gap, indicating a weak SC–coupling between PbTe and Pb. We predict a proximity–induced superconducting gap in PbTe and a softening of Pb SC gap (with respect to bulk). We show that the superconducting anomalous charge density ($\chi(\mathbf{r})$) emerging at the interface is anisotropic. Our findings are robust against lattice deformations (strain) and applied electric field. In particular, the induced SC gap at the interface is resilient under lattice strain.
Figure 1: (a) Side view of the relaxed (4, -5) PbTe/Pb heterostructure stacked along the [001] direction. (b) Average electrostatic potential ($\Delta V_{ave}$) across the interface with a schematic of the energy alignment between a semiconductor and a metal. (c) Schematic energy alignment between a semiconductor and a superconductor, representative of PbTe/Pb. The optimal alignment for the emergence of Majorana Zero Modes occurs when the semiconductor gap and the superconductor gap align in energy.
2 PbTe/Pb Heterostructure: Normal state
[…]
In the normal state, the bands (Figure 2a) cross the Fermi level, resulting in a metallic heterostructure with visible band splitting (mostly near the $M$ point) due to hybridization between PbTe and Pb states. To prove this, we remove the interaction between PbTe and Pb by setting their coupling to zero in the Kohn Sham Hamiltonian. The resulting band structure (Figure 2b) does not exhibit any splitting. We highlight the PbTe contribution to the electronic structure with dark red points in (Figure 2b), proving that the states near the Fermi energy predominantly belong to the semiconductor.
Figure 2: (a) PbTe/Pb normal state band structure. The bands cross the Fermi level, resulting in a metallic heterostructure. (b) Normal state bands with the coupling between the PbTe and Pb sides of the heterostructure set to zero. The color intensity represents the contribution from the PbTe orbitals.
The interaction between the two materials leads to variation in chemical bonds and relative atomic positions near the interface, along with charge transfer from PbTe to Pb. Using Mulliken charge analysis 15, we compute the variation of charge $\Delta Q$ on PbTe ($Q_{PbTe}$) and Pb ($Q_{Pb}$) in the HS with respect to their freestanding nominal values. We find a depletion of charge on PbTe indicating an electron transfer from the PbTe to the Pb side of the HS (Supporting Information, FigureS11).
When a metal comes in contact with a semiconductor, their energy level align, creating a potential energy barrier (Schottky barrier) at the interface due to the difference in their work functions (Figure 1b). We calculate from first-principles the average electrostatic potential ($\Delta V_{ave}(z)$)16 , namely the spatial average of all electrostatic interactions in the system. The average is performed layer-by-layer, and it quantifies the band alignment in heterostructure. It is desirable to have a $\Delta V_{ave}\sim 0$ eV so that the electronic states participating in the superconducting proximity effect are close to the SOC splitted bands of PbTe at $L$ (Figure 1d) and Supporting Information, Figure S8]. In this case, the requirements for emergence of MZMs would be satisfied 7. We predict, instead, a significant difference in the average electrostatic potential across the interface $\Delta V_{ave}\sim 1.2$ eV, with the PbTe side being higher in energy than the Pb side (Figure 1b).
We studied the robustness of these findings against strain and the presence of an external electric field. Our results indicate that $\Delta V_{ave}$ and the normal state electronic DOS near the Fermi level are not significantly affected by varying strain (Table 1 and Figure 3a). We compute the excess charge $\Delta Q$ and observe that it slightly increases by increasing strain on PbTe (Table 3), substantiating the conclusion that the electrostatic environment close to the interface does not change under lattice deformations. Furthermore, the shape and magnitude of the DOS is not affected by the electric field (Figure 3b). The magnitude of the applied electric fields are all considerably above the electrical breakdown point of bulk PbTe which can be estimated to be $6.4\times 10^{-6}$ V/Å 17, while its sign and direction are chosen to offset the difference in electrostatic potential at the interface $\Delta V_{ave}$.
Figure 3: PbTe/Pb normal state DOS for different (a) strain and (b) applied electric field across the interface, represented with different colors.
Table 1: Difference in average electrostatic potential $\Delta V_{ave}$ across the interface of the PbTe/Pb HS for different strain. The (4, -5) HS is highlighted in bold.
strain (PbTe/Pb) [%] $\Delta V_{ave}$ [eV]
1.0%/-7.7% 1.36
2.0%/-6.8% 0.84
3.0%/-5.9% 0.93
4.0%/-5.0% 1.20
5.0%/-4.0% 1.08
6.0%/-3.2% 1.10
7.0%/-2.3% 1.17
8.0%/-1.3% 1.22
9.0%/-0.4% 1.30
3 PbTe/Pb heterostructure: Superconducting state
Figure 4: Superconducting DOS (SC–DOS, $\rho(\varepsilon)$, black continuous line) and anomalous DOS (ADOS, $\chi(\varepsilon)$, dashed green line) for the (4, -5) PbTe/Pb HS. Two distinct energy ranges are resolved within the soft BCS–like SC gap: the difference between the two main coherence peaks $\Delta$ and the induced SC–gap $\Delta_{ind}$, where the SC–DOS vanishes. The dashed black line serves as a reference for zero DOS. (a) SC–DOS projected on individual PbTe (b) and Pb (c) layers, and for the PbTe/Pb HS for diffrent strain values (d).
We used the SIESTA-BdG method to compute the superconducting DOS of bulk Pb (Supporting Information) and PbTe/Pb heterostructures. We demonstrate that a proper initialization of the superconducting pairing potential is essential to avoid unphysical results (Supporting Information Figure S10). Moreover, we demonstrate how the SIESTA-BdG method can be used to investigate various pairing mechanisms and their effects on the superconducting density of states (SC–DOS). Computational details and geometrical aspects of the HS are presented in the Supporting Information Section S6.
Bulk Pb exhibits a hard U–shaped gap of $\Delta\sim 3.0$ meV. The HS, instead, exhibits a soft U–shaped gap with coherence peaks $\Delta=2.0$ meV far apart and an induced gap $\Delta_{ind}=0.6$ meV (Figure 4a). We find that the anomalous density of states (ADOS, $\chi(\varepsilon)$) is dominated by the singlet components (Cooper pair with total spin zero), while the triplet components (Cooper pair with total spin one) are negligible. The shape of the total DOS is determined by the induced gap in the PbTe side of the heterostucture, as indicated by the projected density of states (Figure 4b). The low energy states inside the SC gap are predominantly of PbTe character, hence the HS is in the weak coupling regime 10. The hybridization between Pb and PbTe “poisons” the SC gap, leading to a soft BCS–like superconducting density of states (SC–DOS $\rho(\varepsilon)$). The Pb side of the heterostructure retains its U-shape but the coherence peaks are less sharp and smaller in magnitude than bulk Pb[Figure 4.(c)]. In our simulations, the soft–BCS gap at the SM–SC interface is due to structural relaxation of the atomic structure and states hybridization, as we do not include magnetic, thermal, or dissipative broadening effects 18. We find that strain only slightly affects the shape of the SC DOS in the HS (Figure 4d).
Figure 5: (a) Anomalous density of states $\chi(\mathbf{r})$ along the z-axis in real space. Vertical dashed lines represent the layer position in the PbTe (olive) and Pb (black) regions of the heterostructure. (b) 3D visualization of the anomalous DOS near the interface region. The isosurface level, indicate regions where $\chi(\mathbf{r})$ is constant. To enhance visualization, the isosurface on the PbTe side of the heterostructure (blue) is set to a value 10 times smaller than the Pb side (red).
From the computed anomalous charge density $\chi(\mathbf{r})$, we explain the SC–gap dependence from the position along the heterostructure. On the Pb side, $\chi(\mathbf{r})$ oscillates around a constant value with maxima at the Pb atomic positions (Figure 5a). The oscillations are likely due to the finite size of the simulated system. On the PbTe side, $\chi(\mathbf{r})$ decays from the interface towards the surface with a superconducting coherence length $\xi\sim 14$. The coherence length ($\xi$) was computed by fitting an exponential function to $\chi(\mathbf{r})$.
We visualize the anomalous charge density in (Figure 5b), where the isosurface level indicate regions with constant $\chi(\mathbf{r})$. To enhance visualization, the isosurface on the PbTe side of the heterostructure [blue regions in Figure 5.(b)] is set to a value 10 times smaller. The anomalous charge density penetrates in the PbTe region, it is enhanced at the interface, and it is highly anisotropic. On the Pb side of the HS, the anomalous density is homogenous in the central region (layers 21–22) similar to bulk Pb, while it deforms around the interface and the vacuum region. The anomalous charge density is directly proportional to the superconducting pairing potential 13. Therefore, the pairing potential in the HS is not isotropic. The anisotropy of the pairing potential constitute a signature of unconventional superconductivity.
4 Conclusion
We used a combined Kohn–Sham and Bogoliubov–de Gennes approach to investigate the semiconductor–superconductor (SM–SC) heterostructure composed of PbTe and Pb, focusing on the interfacial properties and the potential applicability of this system for detecting Majorana Zero Modes (MZMs) 19, 9, 7. We discussed the (normal state) metallization of the heterostructure driven by strong hybridization between PbTe and Pb states at the interface. We predict a significant difference in electrostatic potential at the interface, which poses a challenge in achieving the necessary energy alignment between the SM Zeeman gap and SC gap. These findings remain robust against strain and electric field.
In the superconducting state, a soft BCS–like superconducting gap is computed, indicating a weak superconducting coupling between the two systems 10. Moreover, we predicted proximity–induced superconductivity in the PbTe region of the heterostructure. We resolved the superconducting density of states layer by layer, revealing that the induced SC gap is enhanced near the interface, with a softened SC gap on the Pb side. We analyzed the shape and decay of the anomalous charge density $\chi(\mathbf{r})$, revealing an anisotropic pairing potential. We observed a weak superconducting coupling regime at the interface, which prevents the superconductor to overscreen the HS properties. The weak regime favours the device’s tunability with gate and bias voltages, creating a route to get the desirable energy alignment between the two materials. Our findings strengthen the understanding of the physics governing PbTe/Pb hybrid devices and their potential application in hosting Majorana Zero Modes (MZMs).
Acknowledgments
The authors acknowledge the fruitful discussion with Erik Bakkers, Arnold H. Kole and Nils Wittemeier. ZZ acknowledges the research program “Materials for the Quantum Age” (QuMat) for financial support. This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW). RR and ZZ acknowledge financial support from Sector Plan Program 2019-2023. This work was sponsored by NWO-Domain Science for the use of supercomputer facilities. We also acknowledge that the results of this research have been achieved using the Tier-0 PRACE Research Infrastructure resource Discoverer based in Sofia, Bulgaria (OptoSpin project id. 2020225411). This project has received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No 101130384 (QUONDENSATE).
Supporting Information for:
Ab initio study of Proximity-Induced Superconductivity in PbTe/Pb heterostructures
R. Reho, Andrés R. Botello-Méndez, and Zeila Zanolli
The PDF file includes:
Normal and superconducting properties of bulk Pb and PbTe, theoretical methods, details concerning the initialization of the superconducting pairing potential, geometrical aspects of the heterostructures, and supplementary figures.
S5 Supplementary note 1: Computational methods
To model the PbTe/Pb interface, we computed its electronic structure using density functional theory (DFT) as implemented in the SIESTA code 20, 14, including spin-orbit coupling and the PBE exchange-correlation functional 21. We employed optimized norm-conserving Vanderbilt pseudopotentials (ONCVPSP 22) in PSML format 23 from PseudoDojo database 24. The Kohn-Sham equations were solved using standard double-zeta polarized basis sets of localized atomic orbitals on a $60\times 60\times 1$ Monkhorst-Pack grid. Real space integral were evaluated on grids with a cut-off energy of 2000 Ry for the SC state calculations and 500  Ry for the normal state ones. The Fermi-Dirac occupation function was smoothed using an electronic temperature of 0.1 meV, to ensure that the broadening of the electronic states is well below the size of the superconducting gap. All crystal structures were relaxed using the FIRE algorithm with a force threshold of 0.005 eV/Å and a maximum stress tolerance of 0.5 meV/Å3.
The superconducting properties of PbTe/Pb were modeled using the SIESTA–BdG method 13. In this approach a semi-empirical superconducting pairing potential is introduced on top of the self-consistent Kohn-Sham Hamiltonian describing the normal state. We used the fixed–$\Delta$ method with the superconducting pairing potential initialized in real space as a touching spherical hardwells with radius $\mathbf{r}=1.58$ and strength $\bar{\Delta}=1.50$ meV around Pb atoms. We disregard the singlet–triplet mixing effect 25 as the computed contributions from triplet states are considerably lower in magnitude than the singlet ones.
Figure S6: Orbital (a) and spin (b) projected bands structure of bulk Pb in the normal state. In (a) we observe that the electronic states crossing the Fermi level are $p$ and $d$ orbitals. In (b) we observe a spin flip at $W$ for $S_{x}$ and $S_{y}$ spin orientations.
S6 Supplementary note 2: Geometrical aspect of the HS
We constructed the PbTe/Pb HS, stacking 15 layers slab of PbTe cell and 11 layers of Pb along the [001] direction. The HS extends infinitely in the x and y directions. Periodic replicae are separated by 26 Å  vacuum along the z–axis. The PbTe slab was constructed by rotating the in-plane conventional cell by 45 degrees and reconstructing its primitive cell. The Pb slab was constructed by repeating the conventional cell along the [001] direction 12 times and removing one layer to achieve 11 layers of Pb. Farey sequences 26 were employed to identify commensurate cells for the full heterostructure. This method inherently introduces strain in the system. We considered several strained configurations of the PbTe/Pb heterostructure. The strain on each side of the heterostructure is defined as:
$\text{strain}=\frac{\mathrm{a}_{\mathrm{hs}}-\mathrm{a}_{\mathrm{rlx,i}}}{\mathrm{a}_{\mathrm{rlx,i}}}\quad\mathrm{i=Pb,PbTe}$ (S1)
where $\mathrm{a}_{\mathrm{hs}}$ is the lattice parameter of the heterostructure while $\mathrm{a}_{\mathrm{rlx,i}}$ is the relaxed lattice parameter of the constituent materials.
S7 Supplementary note 3: Bulk Pb and bulk PbTe
We present the crystal structure and electronic properties of bulk Pb and PbTe, emphasizing the orbitals’ contribution to the electronic states near the Fermi level. The interaction and hybridization between these states are critical for the proximity–induced superconductivity effect at the PbTe/Pb interface.
Figure S7: (a) SC–DOS (continous lines) and ADOS (dashed lines) of bulk Pb computed with the solution methods: non SCF-BdG (black), Fixed-$\Delta$ (blue), and Full SCF-BdG (red). The Full SCF-BdG method reveals a small splitting of the main coherence peaks. (b) Orbital contribution to the DOS using the Full SCF-BdG method, showing that the main contributions from $p$ and $d$ orbitals, as indicated by the orbitals crossing the Fermi level. Similar results are found using the other two solution methods.
S7.1 Bulk Pb
[…]
We modeled the superconducting state using the SIESTA-BdG code 13, employing three solution methods: non SCF-BdG, Fixed-$\Delta$, and full SCF-BdG. All three solution methods predict a conventional U-shaped SC gap of $\sim$3 meV (Figure S7). We initialize the pairing potential in the superconducting strength representation (on a real-space grid) as touching spherical hardwells with radius $\mathbf{r}=1.58$. In the non SCF–BdG and fixed–$\Delta$ methods, the initial strength of the superconducting pairing potential is $\bar{\Delta}=1.50$ meV. In the full SCF-BdG method, instead, the initial value of the superconducting coupling is $\lambda(\mathbf{r})=80.3$ eV. We computed the orbital projected superconducting density of states (Figure S7) and conclude that Pb $p$ and $d$ are the dominant electronic states close to the Fermi level, and hence relevant for the superconducting properties.
S7.2 Bulk PbTe
[…]
Table 2: Electronic gap ($\Delta E_{L}$) of PbTe as a function of its in-plane lattice parameter. The configuration that matches the value of the lattice parameter for the (4,-5) PbTe/Pb HS is highlighted in bold.
Lattice parameter [Å] $\Delta E_{L}$ [eV]
4.39 0.251
4.44 0.224
4.49 0.147
4.53 0.069
4.58 0.023
4.63 0.058
4.67 0.107
4.72 0.149
4.76 0.181
Figure S8: Electronic band structures for PbTe bulk. (a) Orbital projected band structure. The inset shows an energy region close to the Fermi level. (b) Spin projected band structure. Positive and negative spins are depicted in red and blue, respectively.
Figure S9: Projected DOS of bulk PbTe. Near the Fermi level, the valence band is predominantly influenced by Te $p$ orbitals, whereas the conduction band shows a substantial contribution from Pb $p$ orbitals
S8 Supplementary note 4: Initialization of the Pairing Potential and Its Impact on the Superconducting Properties
In the SIESTA-BdG method one can employ different methods for solving the BdG equations and different approaches for initializing the pairing potential $\Delta$ 13. We tested multiple initializations of the pairing potential. In the main text, we report results obtained within the superconducting strength representation for all simulations.
We compared the superconducting strength and orbital representation (intra–orbital coupling or full–overlap) for the PbTe/Pb heterostructure (Figure S10). The intra–orbital coupling lacks the interactions between different orbitals that can be captured by the superconducting strength representation, making it more localized. As a result, the intra–orbital coupling misses the hybridization and coupling between states belonging to the PbTe and Pb sides of the heterostructure, leading to a U–shaped DOS. However, when we excessively couple states belonging to different orbitals and atoms (full–overlap), the superconducting gap closes. Therefore, initializing the pairing potential in real–space as touching sphere hardwells is a physically intuitive and reasonable approach.
Figure S10: The SC–DOS, $\rho(\varepsilon))$, of the PbTe/Pb heterostructure is computed using the superconducting strength representation method (solid black line) and the orbital representation method. In the orbital representation, we evaluate intra-orbital coupling (solid orange line) and full-overlap coupling (dashed blue line). When pairing occurs between electrons and holes within the same orbital on the same atom (intra–orbital), a conventional U-shaped gap emerges. In contrast, when superconducting coupling involves interactions between different orbitals across different atoms (superconducting strength representation or full–overlap), the hard superconducting gap becomes softened.
S9 Supplementary note 5: Mulliken charge analysis
Table 3: Excess of charge $\Delta Q$ computed for different strain. Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer. We compare the excess of charge on the PbTe side ($\Delta Q_{PbTe}$) of the heterostructure, with the one on the Pb side ($\Delta Q_{Pb}$). The (4, -5) HS is highlighted in bold.
strain (PbTe/Pb) [%] $\Delta Q_{PbTe}$ [1/e] $\Delta Q_{Pb}$ [1/e]
1.0%/-7.7% -0.096 0.096
2.0%/-6.8% -0.098 0.098
3.0%/-5.9% 0.101 0.101
4.0%/-5.0% -0.103 0.103
5.0%/-4.0% -0.107 0.107
6.0%/-3.2% -0.106 0.106
7.0%/-2.3% -0.109 0.109
8.0%/-1.3% -0.114 0.114
9.0%/-0.4% -0.119 0.119
Figure S11: Mulliken charge analysis for PbTe/Pb. Layer-by-layer excess charge. Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer. The red dashed line denote the separation between the PbTe and Pb side of the heterostructure. We notice that at the interface there is a depletion of charge on the PbTe side of the heterostructure, while the opposite is true on the Pb side. The sum of the excess charge on the PbTe and Pb side $\Delta Q$ is reported in the plot.

### v2

@@
-Ab initio study of Proximity-Induced Superconductivity in PbTe/Pb heterostructures
+Proximity-Induced Superconductivity in PbTe/Pb heterostructures from first–principles
 Abstract
-Semiconductor–superconductor hybrid devices have been proposed as promising platforms for detecting and analyzing Majorana zero modes, which find applications in topological quantum computing.
+Semiconductor–superconductor interfaces play a crucial role in various applications, including hybrid circuits, thermometry, devices (bolometers, electronic coolers), detectors for high-energy particle physics, and quantum computing as potential hosts of topologically protected zero-energy modes.
 In this work, we solve the Kohn-Sham Density Functional Theory and Bogoliubov–de Gennes equations to describe the normal and superconducting properties of a PbTe/Pb heterostructure.
-We resolve a proximity–induced superconducting gap on the PbTe side.
-The hybridization between PbTe and Pb causes the emergence of a soft Bardeen–Cooper–Schrieffer–like superconducting gap.
-We compute the anomalous charge density in real space, estimating its decay length and showing that the pairing potential is anisotropic, which is a necessary condition for unconventional superconductivity.
-Contrary to the models that predict Majorana zero modes in these interfaces, we find a significantly large Schottky barrier in the normal state preventing the emergence of zero modes.
-Our findings strengthen the understanding of the physics governing PbTe/Pb hybrid devices and their viability for Majorana zero modes applications.
-keywords:
-superconductivity, semiconductor-superconductor heterostructure, BdG equations
-1 Introduction
-PbTe/Pb nanowires are a promising material platform to host Majorana Zero Modes (MZMs) due to their resilience to charged impurities, high electron mobility, and excellent crystalline quality  1, 2.
-The intrinsic properties of PbTe (narrow bandgap and large Landé g-factor) combined with the superconducting gap ($\Delta\sim$ 3meV) and relatively high critical temperature ($T_{C}\sim$ 7K) of Pb, make these materials advantageous compared to conventional InAs-Al and InSb-Al nanowires 3, 4.
-To host MZMs, a semiconductor–superconductor (SM–SC) interface must meet several key requirements 5, 6: (i) strong spin-orbit coupling and significant Zeeman energy gap ($V_{Z}$) in the semiconductor, (ii) $s$–wave superconductor, (iii) proximity induced superconductivity, and (iv) the heterostructure chemical potential must lie within the SM’s Zeeman gap and the SC gap.
-When the applied Zeeman field exceeds the critical value $V_{C}=\sqrt{\Delta^{2}+\mu^{2}}$ the system transitions into a topological phase, hosting MZMs at its ends 7, 8, 9.
-Cole et al.
-10 developed a model to describe the density of states (DOS) in a generic nanowire and investigated how the SC–coupling strength and the height of the interface barrier affect the shape of the SC gap.
-They identified a hard U–shaped superconducting gap as indicative of a strong SC–coupling regime, where the superconductor overshadows the semiconductor’s electronic properties.
-Conversely, a soft Bardeen–Cooper–Schrieffer–like gap corresponds to a weak SC–coupling regime, where low-energy states of the HS are predominantly of semiconductor character.
-Weak SC–coupling is, therefore, desirable for achieving topological superconductivity 11, 12.
-In this work, we investigate from first–prinicples proximity effects in a PbTe/Pb heterostructure (HS), focusing on the conditions required for MZMs to emerge.
-All our findings are obtained using the recently implemented SIESTA-BdG method 13, 14, which solves the Kohn-Sham Density Functional Theory and Bogoliubov–de Gennes equations to describe on the same footing the normal and superconducting properties.
-We discuss how the electrostatic environment at the interface is critical for the emergence of MZMs.
+We compute the anomalous charge density in real space, estimating its decay length and showing that the pairing potential is anisotropic.
+We demonstrate that superconductivity in the PbTe/Pb interface is resilient against strain.
+In the normal state we find a large Schottky barrier across the interface, resulting in charge transfer from PbTe to Pb.
+We resolve a proximity–induced superconducting gap on the PbTe side, which originates from hybridization between the Pb and PbTe states near the Fermi energy, which occurs in a weak-coupling regime.
+On the Pb side, the superconducting gap appears partially ‘poisoned’, namely less sharp and less wide than bulk Pb.
+Our first-principles simulations provide a quantitative prediction on emergent structural, electronic (charge transfer, potential drops, band offset), and superconducting properties of the PbTe/Pb interface, namely key information for the design of devices where the PbTe/Pb interface plays a central role as, for instance, in core/shell PbTe/Pb nanowires.
+I Introduction
+Several types of nanoscale heterostructures (HS) have emerged as promising building blocks for next-generation quantum technologies.
+These include planar Josephson junctions  [16, 46, 12, 27], topological insulator/superconductor interfaces  [41, 44, 43, 13], and core/shell semiconductor/superconductor (SM/SC) nanowires [18, 23].
+A common feature across these systems is the crucial role of interfacial proximity effects.
+Strain, band alignment, and external fields can significantly modify the dispersion of states near the Fermi level that participate in superconducting coupling, thus influencing the superconducting properties.
+For example, to host Majorana zero modes (MZMs), the interface between SM and ss–wave SC must exhibit  [7, 25]: (i) strong spin-orbit coupling (SOC) and significant Zeeman energy gap ($V_{Z}$) in the SM, (ii) proximity induced superconductivity, and (iii) the heterostructure chemical potential must lie within the SM’s Zeeman gap and the SC gap.
+When the applied Zeeman field exceeds the critical value $V_{C}=\sqrt{\Delta^{2}+\mu^{2}}$ the system transitions into a topological phase, hosting MZMs at its ends [29, 2, 20].
+Experiments confirmed proximity-induced superconductivity in topological insulators (TIs), both in core/shell nanowires [42] and planar junctions [8], with the proximity effect decaying exponentially with the TI thickness.
+PbTe/Pb interfaces are advantageous compared to InAs/Al and InSb/Al ones due to PbTe’s narrow bandgap and large Landé g-factor, combined with Pb’s relatively wide superconducting gap ($\Delta\sim$ 3meV) and critical temperature ($T_{C}\sim$ 7K) [38, 14].
+Theoretical models provide insights into the superconducting properties of both nanowires [18, 23] and planar junctions [12, 27].
+Usually, model results are presented as a phase diagram of (topological) superconductivity in terms of several parameters.
+However, first–principles simulations of Bi2Te3/Nb [32] and Bi2Se3/PdTe [24] emphasize the need of fully including the atomistic details of the interface to accurately compute proximity induced superconductivity.
+In this paper, we investigate from first–principles proximity effects in a PbTe/Pb HS.
+We predict how the structural, electronic (charge transfer, potential drops, band offset), and superconducting properties of the individual materials are modified when they constitute an interface.
+In particular, we find that the properties of the materials are resilient to strain and doping.
+All of our findings are obtained using the recently implemented SIESTA-BdG method [31, 9], which simultaneously solves the Kohn-Sham Density Functional Theory and Bogoliubov–de Gennes equations to describe the normal and superconducting properties on the same footing.
 We find significant band bending at the interface, leading to charge transfer from PbTe to Pb, and affecting the position of the chemical potential.
-In the superconducting state we observe a soft BCS–like SC gap, indicating a weak SC–coupling between PbTe and Pb.
-We predict a proximity–induced superconducting gap in PbTe and a softening of Pb SC gap (with respect to bulk).
-We show that the superconducting anomalous charge density ($\chi(\mathbf{r})$) emerging at the interface is anisotropic.
-Our findings are robust against lattice deformations (strain) and applied electric field.
+In the superconducting state we observe a BCS–like SC gap with a coupling between PbTe and Pb of intermediate strength, namely the states of the HS around the Fermi energy are predominantly of semiconductor character.
+We predict a proximity–induced superconducting gap in PbTe and a poisoning of Pb SC gap (with respect to bulk).
+The relevance of such induced superconductivity extends to physical mechanisms where states near the Fermi surface play a central role.
+Moreover, we show that the superconducting anomalous charge density $\chi(\mathbf{r})$ emerging at the interface is anisotropic in real space.
 In particular, the induced SC gap at the interface is resilient under lattice strain.
-Figure 1: (a) Side view of the relaxed (4, -5) PbTe/Pb heterostructure stacked along the [001] direction.
-(b) Average electrostatic potential ($\Delta V_{ave}$) across the interface with a schematic of the energy alignment between a semiconductor and a metal.
-(c) Schematic energy alignment between a semiconductor and a superconductor, representative of PbTe/Pb.
-The optimal alignment for the emergence of Majorana Zero Modes occurs when the semiconductor gap and the superconductor gap align in energy.
-2 PbTe/Pb Heterostructure: Normal state
-[…]
-In the normal state, the bands (Figure 2a) cross the Fermi level, resulting in a metallic heterostructure with visible band splitting (mostly near the $M$ point) due to hybridization between PbTe and Pb states.
+Figure 1: (a) Side view of the relaxed (4, -5) PbTe/Pb interface stacked along the [001] direction.
+(b) Average electrostatic potential $\Delta V_{ave}$ across the interface, including a schematic of the energy alignment between a semiconductor and a normal metal.
+(c) Schematic band alignment at the interface between a semiconductor and a superconductor.
+II PbTe/Pb Heterostructure: Normal state
+PbTe and Pb {001} planes are not commensurate.
+Hence, in order to build the HS, it is necessary to introduce strain into the unit cell of both materials.
+We employ 15 PbTe and 11 Pb layers, ensuring that the DOS in the central region of PbTe (layers 6–7 Fig 1.(a)] and Pb [layers 21–22 Fig 1.(a)] matches the corresponding bulk values.
+We examined a range of strain from $\sim$1% to $\sim$9% on the PbTe side, and $\sim$ -8% to $\sim$ -0.4% on the Pb side, relaxing the atomic positions for each configuration.
+We denote strain in the HS by a pair of numbers referring to strain on PbTe and Pb, respectively.
+In the following, we focus on the HS with 4% strain on PbTe and -5% strain on Pb, namely (4, -5).
+In the normal state, the bands [Fig.
+2.(a)] cross the Fermi level, resulting in a metallic heterostructure with visible band splitting (mostly near the $M$ point) due to hybridization between PbTe and Pb states, coupling the SC and SM.
 To prove this, we remove the interaction between PbTe and Pb by setting their coupling to zero in the Kohn Sham Hamiltonian.
-The resulting band structure (Figure 2b) does not exhibit any splitting.
-We highlight the PbTe contribution to the electronic structure with dark red points in (Figure 2b), proving that the states near the Fermi energy predominantly belong to the semiconductor.
+The resulting band structure [Fig.
+2.(b)] does not exhibit any splitting.
+We highlight the PbTe contribution to the electronic structure with dark red points in [Fig.
+2.(b)], proving that the states near the Fermi energy predominantly belong to the semiconductor.
 Figure 2: (a) PbTe/Pb normal state band structure.
@@
 The interaction between the two materials leads to variation in chemical bonds and relative atomic positions near the interface, along with charge transfer from PbTe to Pb.
-Using Mulliken charge analysis 15, we compute the variation of charge $\Delta Q$ on PbTe ($Q_{PbTe}$) and Pb ($Q_{Pb}$) in the HS with respect to their freestanding nominal values.
-We find a depletion of charge on PbTe indicating an electron transfer from the PbTe to the Pb side of the HS (Supporting Information, FigureS11).
-When a metal comes in contact with a semiconductor, their energy level align, creating a potential energy barrier (Schottky barrier) at the interface due to the difference in their work functions (Figure 1b).
-We calculate from first-principles the average electrostatic potential ($\Delta V_{ave}(z)$)16 , namely the spatial average of all electrostatic interactions in the system.
-The average is performed layer-by-layer, and it quantifies the band alignment in heterostructure.
-It is desirable to have a $\Delta V_{ave}\sim 0$ eV so that the electronic states participating in the superconducting proximity effect are close to the SOC splitted bands of PbTe at $L$ (Figure 1d) and Supporting Information, Figure S8].
-In this case, the requirements for emergence of MZMs would be satisfied 7.
-We predict, instead, a significant difference in the average electrostatic potential across the interface $\Delta V_{ave}\sim 1.2$ eV, with the PbTe side being higher in energy than the Pb side (Figure 1b).
+Using Mulliken charge analysis [21], we compute the variation of charge $\Delta Q$ on PbTe ($Q_{PbTe}$) and Pb ($Q_{Pb}$) in the HS with respect to their freestanding nominal values.
+We find a depletion of charge on PbTe indicating an electron transfer from the PbTe to the Pb side of the HS (Table D1 and Fig.D6).
+The presence of PbTe states crossing the Fermi level and the moderate charge transfer towards Pb, indicate an intermediate-strength coupling regime between the two materials.
+When a semiconductor and a metal form an interface, their energy band edges align, creating a potential energy barrier (Schottky barrier) due to the difference in their work functions [Fig.
+1(b)].
+We calculate from first-principles the average electrostatic potential $\Delta V_{ave}(z)$  [4], namely the spatial average of all electrostatic interactions in the system.
+The average is performed layer-by-layer, and it quantifies the band alignment in heterostructures.
+We predict a significant difference in the average electrostatic potential across the interface $\Delta V_{ave}\sim 1.2$ eV, with the PbTe side being higher in energy than the Pb side [Fig.
+1(b)].
 We studied the robustness of these findings against strain and the presence of an external electric field.
-Our results indicate that $\Delta V_{ave}$ and the normal state electronic DOS near the Fermi level are not significantly affected by varying strain (Table 1 and Figure 3a).
-We compute the excess charge $\Delta Q$ and observe that it slightly increases by increasing strain on PbTe (Table 3), substantiating the conclusion that the electrostatic environment close to the interface does not change under lattice deformations.
-Furthermore, the shape and magnitude of the DOS is not affected by the electric field (Figure 3b).
-The magnitude of the applied electric fields are all considerably above the electrical breakdown point of bulk PbTe which can be estimated to be $6.4\times 10^{-6}$ V/Å 17, while its sign and direction are chosen to offset the difference in electrostatic potential at the interface $\Delta V_{ave}$.
+Our results indicate that $\Delta V_{ave}$ and the normal state electronic DOS near the Fermi level are not significantly affected by varying strain [Table 1 and Fig.
+3.(a)].
+We compute the excess charge $\Delta Q$ and observe that it slightly increases by increasing strain on PbTe (Table D1), substantiating the conclusion that the electrostatic environment close to the interface does not change under lattice deformations.
+Furthermore, the shape and magnitude of the DOS is not affected by the electric field [Fig.
+3(b)].
+The magnitude of the applied electric fields are all considerably above the electrical breakdown point of bulk PbTe which can be estimated to be $6.4\times 10^{-6}$ V/Å [40], while its sign and direction are chosen to offset the difference in electrostatic potential at the interface $\Delta V_{ave}$.
 Figure 3: PbTe/Pb normal state DOS for different (a) strain and (b) applied electric field across the interface, represented with different colors.
@@
 9.0%/-0.4% 1.30
-3 PbTe/Pb heterostructure: Superconducting state
-Figure 4: Superconducting DOS (SC–DOS, $\rho(\varepsilon)$, black continuous line) and anomalous DOS (ADOS, $\chi(\varepsilon)$, dashed green line) for the (4, -5) PbTe/Pb HS.
-Two distinct energy ranges are resolved within the soft BCS–like SC gap: the difference between the two main coherence peaks $\Delta$ and the induced SC–gap $\Delta_{ind}$, where the SC–DOS vanishes.
-The dashed black line serves as a reference for zero DOS.
-(a) SC–DOS projected on individual PbTe (b) and Pb (c) layers, and for the PbTe/Pb HS for diffrent strain values (d).
-We used the SIESTA-BdG method to compute the superconducting DOS of bulk Pb (Supporting Information) and PbTe/Pb heterostructures.
-We demonstrate that a proper initialization of the superconducting pairing potential is essential to avoid unphysical results (Supporting Information Figure S10).
-Moreover, we demonstrate how the SIESTA-BdG method can be used to investigate various pairing mechanisms and their effects on the superconducting density of states (SC–DOS).
-Computational details and geometrical aspects of the HS are presented in the Supporting Information Section S6.
-Bulk Pb exhibits a hard U–shaped gap of $\Delta\sim 3.0$ meV.
-The HS, instead, exhibits a soft U–shaped gap with coherence peaks $\Delta=2.0$ meV far apart and an induced gap $\Delta_{ind}=0.6$ meV (Figure 4a).
-We find that the anomalous density of states (ADOS, $\chi(\varepsilon)$) is dominated by the singlet components (Cooper pair with total spin zero), while the triplet components (Cooper pair with total spin one) are negligible.
-The shape of the total DOS is determined by the induced gap in the PbTe side of the heterostucture, as indicated by the projected density of states (Figure 4b).
-The low energy states inside the SC gap are predominantly of PbTe character, hence the HS is in the weak coupling regime 10.
-The hybridization between Pb and PbTe “poisons” the SC gap, leading to a soft BCS–like superconducting density of states (SC–DOS $\rho(\varepsilon)$).
-The Pb side of the heterostructure retains its U-shape but the coherence peaks are less sharp and smaller in magnitude than bulk Pb[Figure 4.(c)].
-In our simulations, the soft–BCS gap at the SM–SC interface is due to structural relaxation of the atomic structure and states hybridization, as we do not include magnetic, thermal, or dissipative broadening effects 18.
-We find that strain only slightly affects the shape of the SC DOS in the HS (Figure 4d).
+III PbTe/Pb heterostructure: Superconducting state
+Figure 4: (a) Superconducting DOS (SC–DOS, $\rho(\varepsilon)$, black continuous line) and anomalous DOS (ADOS, $\chi(\varepsilon)$, dashed green line) for the (4, -5) PbTe/Pb HS, with coherence peaks at $\pm\Delta/2\sim\pm 1$ eV.
+The dashed black line indicates zero DOS.
+SC–DOS projected on individual (b) PbTe and (c) Pb layers, and (d) for the PbTe/Pb HS for diffrent strain values.
+We used the SIESTA-BdG method to compute the superconducting DOS and band structure of PbTe/Pb heterostructures finding a proximity-induced SC gap, an intermediate-strength coupling between the two materials, and an anisotropic SC gap in real and reciprocal space.
+Computational details and geometrical aspects of the HS are presented in Section IV.
+More than fifty years ago, experimental measures [6, 19] of bulk Pb revealed an anisotropic U–shaped SC-DOS with two superconducting gaps of width $\sim 2.34-2.5$ and $\sim 2.7-2.8$ meV depending on the crystal orientation.
+Each coherence peak is split in two, with an energy separation of $\sim 0.1-0.2$ meV.
+Splitting of the coherence peaks in bulk Pb has been theoretically predicted with SIESTA–BdG [31] and KKR-BdG methods [33], and is further discussed in Appendix 1.
+Anisotropies in the superconducting gap are also visible in the BdG spectrum of the HS (Figure D7).
+Indeed, due to heterostructuring, the electrostatic environment of Pb atoms changes layer-by-layer, leading to multiple superconducting gaps at different k–points in the BdG spectrum.
+The HS SC gap is U–shaped with coherence peaks $\Delta=2.0$ meV far apart [Fig.
+4.(a)], reflecting an intermediate-strength coupling regime.
+The projected DOS [Fig.
+4.(b) and (c)] show a proximity-induced gap on PbTe and a poisoned SC gap on Pb, namely a SC gap which does not feature a sharp U shape and has reduced width with respect to Pb bulk.
+The shape of the total DOS is determined by the induced gap in the PbTe side of the heterostucture, as indicated by the projected density of states [Fig.
+4.(b)].
+The energy states inside the SC gap are predominantly of PbTe character, hence the HS is in the intermediate coupling regime.
+The hybridization between Pb and PbTe “poisons” the SC gap, leading to a BCS–like superconducting density of states (SC–DOS $\rho(\varepsilon)$).
+The Pb side of the heterostructure retains its U-shape but the coherence peaks are less sharp and smaller in magnitude than bulk Pb [Fig.
+4.(c)].
+In our simulations, the poisoned BCS–like gap at the SM–SC interface is due to structural relaxation of the atomic structure and states hybridization, as we do not include magnetic, thermal, or dissipative broadening effects [37].
+We find that strain only slightly affects the shape of the SC–DOS in the HS [Fig.
+4.(d)].
 Figure 5: (a) Anomalous density of states $\chi(\mathbf{r})$ along the z-axis in real space.
@@
 (b) 3D visualization of the anomalous DOS near the interface region.
-The isosurface level, indicate regions where $\chi(\mathbf{r})$ is constant.
+We set the isosurface level to a fixed value ($1.7\times 10^{-7}$ 1/Bohr3), which indicate regions where $\chi(\mathbf{r})$ is constant.
 To enhance visualization, the isosurface on the PbTe side of the heterostructure (blue) is set to a value 10 times smaller than the Pb side (red).
-From the computed anomalous charge density $\chi(\mathbf{r})$, we explain the SC–gap dependence from the position along the heterostructure.
-On the Pb side, $\chi(\mathbf{r})$ oscillates around a constant value with maxima at the Pb atomic positions (Figure 5a).
+The anomalous density of states (ADOS, $\chi(\varepsilon)$) is dominated by the singlet components (Cooper pair with total spin zero), while the triplet components (Cooper pair with total spin one) are negligible both for buk Pb (Fig.
+B2) and the HS (Fig.
+D8).
+We compute the singlet and triplet components following the Balian-Werthamer (BW) [5] representation including SOC.
+Saunderson et al.
+[34] investigated the difference between solving the BdG equations in the collinear or SOC case.
+The authors propose an orbital Rashba Edelstein effect that leads to enhanced triplet densities for bulk Pb in the absence of SOC.
+Notably, a similar weakening of the triplet density in case of SOC is reported here.
+The anomalous charge density $\chi(\mathbf{r})$ exhibit atomic scale variations, which we can investigate as a function of the position along the heterostructure.
+On the Pb side, $\chi(\mathbf{r})$ oscillates around a constant value with maxima at the Pb atomic positions [Fig.
+5.(a)].
 The oscillations are likely due to the finite size of the simulated system.
-On the PbTe side, $\chi(\mathbf{r})$ decays from the interface towards the surface with a superconducting coherence length $\xi\sim 14$.
-The coherence length ($\xi$) was computed by fitting an exponential function to $\chi(\mathbf{r})$.
-We visualize the anomalous charge density in (Figure 5b), where the isosurface level indicate regions with constant $\chi(\mathbf{r})$.
-To enhance visualization, the isosurface on the PbTe side of the heterostructure [blue regions in Figure 5.(b)] is set to a value 10 times smaller.
+On the PbTe side, $\chi(\mathbf{r})$ decays from the interface towards the surface.
+We fitted an exponential function to $\chi(\mathbf{r})$ and estimated its decay length to be $\eta\sim 14$ Å.
+The estimated decay length $\eta$ is not directly related to the superconducting coherence length of Pb, which is estimated to be $\sim 80$ nm [28].
+Here, $\eta$ is a result of an effective pairing potential that describes the average build-up of anomalous charge densities in the superconductor.
+Finite size effects on superconducting properties of Pb nanofilms [35, 47] have been explained in terms of quantum well states crossing the Fermi level and the spacing between adjacent layers.
+In our simulations, the Pb side of the HS is extremely thin compared to the superconducting coherence length of Pb.
+Consequently, the characteristic energy spacing of quantum well states become significantly larger than the superconducting gap.
+Quantum confinement might play an important role in the effective electron phonon coupling, and thus the pairing potential.
+This effect could be included in the initialization of $\Delta$.
+However, since here we are interested in the interfacial effects, we neglect any quantum size effects.
+As discussed above, the superconducting gap $\Delta(\mathbf{k})$ exhibits anisotropy in k-space, meaning its magnitude varies across different k-points.
+The anomalous charge density is directly related to the pairing potential by the following expression:
+$\Delta_{\mu\nu}^{\alpha\beta}(\mathbf{k})=$ $-\sum_{l=0,1,2,3}\sum_{\gamma}i\sigma_{l}^{\alpha\gamma}\sigma_{2}^{\gamma\beta}\times$ (1)
+$\int_{\mathrm{BZ}}\mathrm{~d}\mathbf{k}^{\prime}\lambda_{l}^{\mu\nu}\left(\mathbf{k}-\mathbf{k}^{\prime}\right)\chi_{\mu\nu}^{l}\left(\mathbf{k}^{\prime}\right),$
+where $\mu$ and $\nu$ label the basis orbitals, $\alpha$ and $\beta$ are spin indices, $\sigma_{l}$ with $l=0,1,2,3$ are the Pauli matrices, and $\lambda^{\mu\nu}$ are the orbital projected superconducting couplings [31].
+We visualize the anomalous charge density in [Fig.
+5.(b)], where the isosurface level indicate regions with constant $\chi(\mathbf{r})$.
+To enhance visualization, the isosurface on the PbTe side of the heterostructure [blue regions in Fig.
+5.(b)] is set to a value 10 times smaller.
 The anomalous charge density penetrates in the PbTe region, it is enhanced at the interface, and it is highly anisotropic.
 On the Pb side of the HS, the anomalous density is homogenous in the central region (layers 21–22) similar to bulk Pb, while it deforms around the interface and the vacuum region.
-The anomalous charge density is directly proportional to the superconducting pairing potential 13.
 Therefore, the pairing potential in the HS is not isotropic.
-The anisotropy of the pairing potential constitute a signature of unconventional superconductivity.
-4 Conclusion
-We used a combined Kohn–Sham and Bogoliubov–de Gennes approach to investigate the semiconductor–superconductor (SM–SC) heterostructure composed of PbTe and Pb, focusing on the interfacial properties and the potential applicability of this system for detecting Majorana Zero Modes (MZMs) 19, 9, 7.
+IV Computational methods
+To model the PbTe/Pb interface, we computed its electronic structure using density functional theory (DFT) as implemented in the SIESTA code [36, 9], including spin-orbit coupling and the PBE exchange-correlation functional [26].
+We employed optimized norm-conserving Vanderbilt pseudopotentials (ONCVPSP [11]) in PSML format [10] from PseudoDojo database [39].
+The Kohn-Sham equations were solved using standard double-zeta polarized basis sets of localized atomic orbitals on a $60\times 60\times 1$ Monkhorst-Pack grid.
+Real space integral were evaluated on grids with a cut-off energy of 2000 Ry for the SC state calculations and 500  Ry for the normal state ones.
+The Fermi-Dirac occupation function was smoothed using an electronic temperature of 0.1 meV, to ensure that the broadening of the electronic states is well below the size of the superconducting gap.
+All crystal structures were relaxed using the FIRE algorithm with a force threshold of 0.005 eV/Å and a maximum stress tolerance of 0.5 meV/Å3.
+[…]
+Our approach utilizes an effective pairing potential that describes the average build-up of anomalous charge densities in the superconductor, providing a robust framework for interpreting the physical phenomena.
+Therefore, we do not expect large variations on the superconducting properties of the HS as the Pb thickness changes, even though finite-size effects can significantly influence the SC-DOS in planar Josephson junctions.
+[1, 45, 35] We demonstrate that a proper initialization of the superconducting pairing potential is essential to avoid unphysical results (Appendix C, Fig.
+C5).
+Moreover, we demonstrate how the SIESTA-BdG method can be used to investigate various pairing mechanisms and their effects on the superconducting density of states (SC–DOS).
+V Conclusion
+We used a combined Kohn–Sham and Bogoliubov–de Gennes approach to investigate the semiconductor–superconductor (SM–SC) heterostructure composed of PbTe and Pb, focusing on the interfacial properties.
 We discussed the (normal state) metallization of the heterostructure driven by strong hybridization between PbTe and Pb states at the interface.
@@
 These findings remain robust against strain and electric field.
-In the superconducting state, a soft BCS–like superconducting gap is computed, indicating a weak superconducting coupling between the two systems 10.
-Moreover, we predicted proximity–induced superconductivity in the PbTe region of the heterostructure.
-We resolved the superconducting density of states layer by layer, revealing that the induced SC gap is enhanced near the interface, with a softened SC gap on the Pb side.
+In the superconducting state, a BCS–like superconducting gap is computed, indicating a intermediate superconducting coupling between the two systems.
+Moreover, we predicted proximity–induced superconductivity in the PbTe region of the HS.
+We resolved the superconducting density of states layer by layer, revealing that the induced SC gap is enhanced near the interface, with a poisoned SC gap on the Pb side.
 We analyzed the shape and decay of the anomalous charge density $\chi(\mathbf{r})$, revealing an anisotropic pairing potential.
@@
 The weak regime favours the device’s tunability with gate and bias voltages, creating a route to get the desirable energy alignment between the two materials.
-Our findings strengthen the understanding of the physics governing PbTe/Pb hybrid devices and their potential application in hosting Majorana Zero Modes (MZMs).
+The field of topological quantum materials is increasingly moving towards identifying optimal design parameters to control and enhance the functionality of hybrid devices.
+Our results naturally fit in this direction, by providing first-principles quantitative predictions of the superconducting and electronic properties of the PbTe/Pb semiconductor/superconductor interface and their resiliance to strain and electric field.
+We expect our findings to benefit both experimentalists and theoreticians alike, due to the key role played by this interface in various technologically relevant devices, among which core/shell semiconductor/superconductor nanowires.
 Acknowledgments
@@
 This work was sponsored by NWO-Domain Science for the use of supercomputer facilities.
+This publication is part of the project ”Quantum Materials by Design” with file number 2024.012 of the research programme Computing Time on National Computer Facilities which is (partly) financed by the Dutch Research Council (NWO) under the grant https://doi.org/10.61686/JUXDK41645 .
 We also acknowledge that the results of this research have been achieved using the Tier-0 PRACE Research Infrastructure resource Discoverer based in Sofia, Bulgaria (OptoSpin project id.
@@
 This project has received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No 101130384 (QUONDENSATE).
-Supporting Information for:
-Ab initio study of Proximity-Induced Superconductivity in PbTe/Pb heterostructures
-R. Reho, Andrés R. Botello-Méndez, and Zeila Zanolli
-The PDF file includes:
-Normal and superconducting properties of bulk Pb and PbTe, theoretical methods, details concerning the initialization of the superconducting pairing potential, geometrical aspects of the heterostructures, and supplementary figures.
-S5 Supplementary note 1: Computational methods
-To model the PbTe/Pb interface, we computed its electronic structure using density functional theory (DFT) as implemented in the SIESTA code 20, 14, including spin-orbit coupling and the PBE exchange-correlation functional 21.
-We employed optimized norm-conserving Vanderbilt pseudopotentials (ONCVPSP 22) in PSML format 23 from PseudoDojo database 24.
-The Kohn-Sham equations were solved using standard double-zeta polarized basis sets of localized atomic orbitals on a $60\times 60\times 1$ Monkhorst-Pack grid.
-Real space integral were evaluated on grids with a cut-off energy of 2000 Ry for the SC state calculations and 500  Ry for the normal state ones.
-The Fermi-Dirac occupation function was smoothed using an electronic temperature of 0.1 meV, to ensure that the broadening of the electronic states is well below the size of the superconducting gap.
-All crystal structures were relaxed using the FIRE algorithm with a force threshold of 0.005 eV/Å and a maximum stress tolerance of 0.5 meV/Å3.
-The superconducting properties of PbTe/Pb were modeled using the SIESTA–BdG method 13.
-In this approach a semi-empirical superconducting pairing potential is introduced on top of the self-consistent Kohn-Sham Hamiltonian describing the normal state.
-We used the fixed–$\Delta$ method with the superconducting pairing potential initialized in real space as a touching spherical hardwells with radius $\mathbf{r}=1.58$ and strength $\bar{\Delta}=1.50$ meV around Pb atoms.
-We disregard the singlet–triplet mixing effect 25 as the computed contributions from triplet states are considerably lower in magnitude than the singlet ones.
-Figure S6: Orbital (a) and spin (b) projected bands structure of bulk Pb in the normal state.
-In (a) we observe that the electronic states crossing the Fermi level are $p$ and $d$ orbitals.
-In (b) we observe a spin flip at $W$ for $S_{x}$ and $S_{y}$ spin orientations.
-S6 Supplementary note 2: Geometrical aspect of the HS
-We constructed the PbTe/Pb HS, stacking 15 layers slab of PbTe cell and 11 layers of Pb along the [001] direction.
+VI Data availability
+The data of this work are available at the following NOMAD repository [30].
+VII keywords
+proximity interaction, superconductivity, semiconductor-superconductor heterostructures, first-principles simulations, Bogoliubov de Gennes equations
+Appendix A Geometrical aspect of the HS
+We constructed the PbTe/Pb HS by stacking 15 layers of PbTe and 11 layers of Pb along the [001] direction.
 The HS extends infinitely in the x and y directions.
-Periodic replicae are separated by 26 Å  vacuum along the z–axis.
+Periodic replicae are separated by 26 Å of vacuum along the z–axis.
 The PbTe slab was constructed by rotating the in-plane conventional cell by 45 degrees and reconstructing its primitive cell.
-The Pb slab was constructed by repeating the conventional cell along the [001] direction 12 times and removing one layer to achieve 11 layers of Pb.
-Farey sequences 26 were employed to identify commensurate cells for the full heterostructure.
+Farey sequences [22] were employed to identify commensurate cells for the full heterostructure.
 This method inherently introduces strain in the system.
@@
 The strain on each side of the heterostructure is defined as:
-$\text{strain}=\frac{\mathrm{a}_{\mathrm{hs}}-\mathrm{a}_{\mathrm{rlx,i}}}{\mathrm{a}_{\mathrm{rlx,i}}}\quad\mathrm{i=Pb,PbTe}$ (S1)
+$\text{strain}=\frac{\mathrm{a}_{\mathrm{hs}}-\mathrm{a}_{\mathrm{rlx,i}}}{\mathrm{a}_{\mathrm{rlx,i}}}\quad\mathrm{i=Pb,PbTe}$ (A)
 where $\mathrm{a}_{\mathrm{hs}}$ is the lattice parameter of the heterostructure while $\mathrm{a}_{\mathrm{rlx,i}}$ is the relaxed lattice parameter of the constituent materials.
-S7 Supplementary note 3: Bulk Pb and bulk PbTe
+Appendix B Bulk Pb and bulk PbTe
 We present the crystal structure and electronic properties of bulk Pb and PbTe, emphasizing the orbitals’ contribution to the electronic states near the Fermi level.
 The interaction and hybridization between these states are critical for the proximity–induced superconductivity effect at the PbTe/Pb interface.
-Figure S7: (a) SC–DOS (continous lines) and ADOS (dashed lines) of bulk Pb computed with the solution methods: non SCF-BdG (black), Fixed-$\Delta$ (blue), and Full SCF-BdG (red).
-The Full SCF-BdG method reveals a small splitting of the main coherence peaks.
-(b) Orbital contribution to the DOS using the Full SCF-BdG method, showing that the main contributions from $p$ and $d$ orbitals, as indicated by the orbitals crossing the Fermi level.
-Similar results are found using the other two solution methods.
-S7.1 Bulk Pb
+1 Bulk Pb
 […]
-We modeled the superconducting state using the SIESTA-BdG code 13, employing three solution methods: non SCF-BdG, Fixed-$\Delta$, and full SCF-BdG.
-All three solution methods predict a conventional U-shaped SC gap of $\sim$3 meV (Figure S7).
-We initialize the pairing potential in the superconducting strength representation (on a real-space grid) as touching spherical hardwells with radius $\mathbf{r}=1.58$.
-In the non SCF–BdG and fixed–$\Delta$ methods, the initial strength of the superconducting pairing potential is $\bar{\Delta}=1.50$ meV.
-In the full SCF-BdG method, instead, the initial value of the superconducting coupling is $\lambda(\mathbf{r})=80.3$ eV.
-We computed the orbital projected superconducting density of states (Figure S7) and conclude that Pb $p$ and $d$ are the dominant electronic states close to the Fermi level, and hence relevant for the superconducting properties.
-S7.2 Bulk PbTe
+Figure B1: Orbital (a) and spin (b) projected band structure of bulk Pb in the normal state.
+In (a) we observe that the electronic states crossing the Fermi level are $p$ and $d$ orbitals.
+In (b) we observe a spin flip at $W$ for $S_{x}$ and $S_{y}$ spin orientations.
+Figure B2: SC–DOS $\rho(\varepsilon)$ (continous lines) and ADOS $\chi(\varepsilon)$ (dashed lines) for bulk Pb computed with the solution methods (a) non SCF-BdG, (b) fixed-$\Delta$, and (c) full SCF-BdG.
+Panel (d) is the orbital projected normal density of states.
+Some lines at zero DOS are not visible as they are hidden behind by the magenta full line.
+We computed the superconducting state using the SIESTA-BdG code [31], including SOC and employing three solution methods: non SCF-BdG, fixed-$\Delta$, and full SCF-BdG.
+All three solution methods predict a conventional U-shaped SC gap of $\sim$3 meV (Fig.
+B2).
+In the following, we discuss the components of the normal $\rho$ and anomalous $\chi$ densities, highlighting how those are obtained from the projected densities of states $\rho^{\alpha\beta}_{\mu\nu}$ and $\chi^{\alpha\beta}_{\mu\nu}$ computed by the code.
+The projected densities of states are given by [31]
+$\rho_{\mu\nu}^{\alpha\beta}(\mathbf{k})=$ $\sum_{i}f\left(\varepsilon_{i}(\mathbf{k})\right)u_{i\mu}^{*\alpha}(\mathbf{k})u_{i\nu}^{\beta}(\mathbf{k})$ (B)
+$+\sum_{i}\bar{f}\left(\varepsilon_{i}(-\mathbf{k})\right)v_{i\mu}^{\alpha}(-\mathbf{k})v_{i\nu}^{*\beta}(-\mathbf{k}),$
+$\chi_{\mu\nu}^{\alpha\beta}(\mathbf{k})=$ $\sum_{i}f\left(\varepsilon_{i}(\mathbf{k})\right)u_{i\nu}^{\beta}(\mathbf{k})v_{n\mu}^{*\alpha}(\mathbf{k})$
+$+\sum\bar{f}\left(\varepsilon_{i}(-\mathbf{k})\right)u_{i\mu}^{\alpha}(-\mathbf{k})v_{i\nu}^{*\beta}(-\mathbf{k}).$
+The tensor structure describes how the weight is distributed among orbitals and spin degrees of freedom.
+Each projected density of states contains an ‘electron” (proportional to $f(\varepsilon_{i}({\mathbf{k}}))$) and ‘hole” (proportional to $\bar{f}(\varepsilon_{i}({\mathbf{k}}))$) contribution, corresponding to the two components of the BdG quasiparticles.
+The spin sector of $\rho^{\alpha\beta}_{\mu\nu}$ is expanded in the basis of Pauli matrices ($\sigma^{0},\sigma^{x},\sigma^{y},\sigma^{z}$).
+This yields one spin-summed term and a three components term, conventionally labelled as $x,y,z$.
+For the anomalous density $\chi^{\alpha\beta}_{\mu\nu}$ the same decomposition is rearranged into singlet (S) and triplet (T) components according to the Balian-Werthamer representation [5].
+Summing over all orbital and spin indices one recovers the scalar density of states ($\rho(\varepsilon)$ and $\chi(\varepsilon)$).
+In Fig.
+B2 we report the total (spin-summed) normal density as sum of the electron and hole components $\rho_{tot_{e}}+\rho_{tot_{h}}$, its $i=x,y,z$ spin-components $\rho_{i}=\rho_{e_{i}}+\rho_{h_{i}}$, and the singlet and triplet components of the anomalous density $\chi_{S}$ and $\chi_{T}$.
+The z component of the SC density of states $\rho(\varepsilon)_{z}$ exhibits small oscillations around zero when computed using the non-SCF-BdG and fixed-$\Delta$ methods but vanishes when computed using the full SCF-BdG method, which is the most accurate one [Fig.
+B2.(a-c)].
+The x and y components of the normal density are always identically zero, regardless of the solution method employed.
+The triplet components $\chi(\varepsilon)_{T_{x}}$ is almost zero for non SCF-BdG and fixed-$\Delta$ solution methods, and identically zero for the full SCF-BdG one.
+The $\chi(\varepsilon)_{T_{y}}$ and $\chi(\varepsilon)_{T_{z}}$ components of $\chi$ are always zero.
+The small oscillations around zero for $\rho(\varepsilon)$ and $\chi(\varepsilon)$ are due to numerical noise.
+The oscillations appear only for the non SCF-BdG and fixed-$\Delta$ solution methods because in these cases the anomalous components are not updated, namely they are kept constant with respect to the initial assumption.
+This further validates our implementation and explains why the DOS of triplet components of $\chi(\varepsilon)$ are symmetrical with respect to the $E_{F}$ in the non SCF-BdG and fixed- $\Delta$ cases.
 […]
-Table 2: Electronic gap ($\Delta E_{L}$) of PbTe as a function of its in-plane lattice parameter.
+2 Bulk PbTe
+[…]
+Figure B3: Electronic band structures for PbTe bulk.
+(a) Orbital projected band structure.
+The inset shows an energy region close to the Fermi level.
+(b) Spin projected band structure.
+Positive and negative spins are depicted in red and blue, respectively.
+Table B1: Electronic gap ($\Delta E_{L}$) of PbTe as a function of its in-plane lattice parameter.
 The configuration that matches the value of the lattice parameter for the (4,-5) PbTe/Pb HS is highlighted in bold.
@@
 4.76 0.181
-Figure S8: Electronic band structures for PbTe bulk.
-(a) Orbital projected band structure.
-The inset shows an energy region close to the Fermi level.
-(b) Spin projected band structure.
-Positive and negative spins are depicted in red and blue, respectively.
-Figure S9: Projected DOS of bulk PbTe.
+Figure B4: Projected DOS of bulk PbTe.
 Near the Fermi level, the valence band is predominantly influenced by Te $p$ orbitals, whereas the conduction band shows a substantial contribution from Pb $p$ orbitals
-S8 Supplementary note 4: Initialization of the Pairing Potential and Its Impact on the Superconducting Properties
-In the SIESTA-BdG method one can employ different methods for solving the BdG equations and different approaches for initializing the pairing potential $\Delta$ 13.
+Appendix C PbTe/Pb HS: Initialization of the Pairing Potential and Its Impact on the Superconducting Properties
+Figure C5: The SC–DOS, $\rho(\varepsilon))$, of the (4,-5) PbTe/Pb HS is computed using the superconducting strength representation method (solid black line) and the orbital representation method.
+In the orbital representation, we evaluate intra-orbital coupling (solid orange line) and full-overlap coupling (dashed blue line).
+When pairing occurs between electrons and holes within the same orbital on the same atom (intra–orbital), a conventional U-shaped gap emerges.
+In contrast, when superconducting coupling involves interactions between different orbitals across different atoms (superconducting strength representation or full–overlap), the hard superconducting gap becomes poisoned.
+In the SIESTA-BdG method one can employ different methods for solving the BdG equations and different approaches for initializing the pairing potential $\Delta$ [31].
 We tested multiple initializations of the pairing potential.
 In the main text, we report results obtained within the superconducting strength representation for all simulations.
-We compared the superconducting strength and orbital representation (intra–orbital coupling or full–overlap) for the PbTe/Pb heterostructure (Figure S10).
+We compared the superconducting strength and orbital representation (intra–orbital coupling or full–overlap) for the PbTe/Pb heterostructure (Fig.
+C5).
 The intra–orbital coupling lacks the interactions between different orbitals that can be captured by the superconducting strength representation, making it more localized.
@@
 Therefore, initializing the pairing potential in real–space as touching sphere hardwells is a physically intuitive and reasonable approach.
-Figure S10: The SC–DOS, $\rho(\varepsilon))$, of the PbTe/Pb heterostructure is computed using the superconducting strength representation method (solid black line) and the orbital representation method.
-In the orbital representation, we evaluate intra-orbital coupling (solid orange line) and full-overlap coupling (dashed blue line).
-When pairing occurs between electrons and holes within the same orbital on the same atom (intra–orbital), a conventional U-shaped gap emerges.
-In contrast, when superconducting coupling involves interactions between different orbitals across different atoms (superconducting strength representation or full–overlap), the hard superconducting gap becomes softened.
-S9 Supplementary note 5: Mulliken charge analysis
-Table 3: Excess of charge $\Delta Q$ computed for different strain.
+Appendix D Additional figures and tables
+This sections presents additional figures and tables related to the PbTe/Pb HS.
+We performed a Mulliken charge analysis and calculated the excess charge $\Delta Q$ under different strain conditions, finding that $\Delta Q$ remains robust agains strain variations.
+(Table D1).
+A layer-by-layer decomposition of $\Delta Q$ (Fig.
+D6) reveals charge depletion on the PbTe side of the interface, with a corresponding accumulation on the Pb side.
+Fig.
+D7 shows the BdG spectrum of the (4,-5) PbTe/Pb HS.
+The SC gap $\Delta(\mathbf{k})$ is anisotropic, as evidenced by the variations in the width of the splitting of the BdG dispersion along the Brillouin Zone
+Figure D6: Mulliken charge analysis for (4,-5) PbTe/Pb HS.
+Layer-by-layer excess charge.
+Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer.
+The red dashed line denote the separation between the PbTe and Pb side of the heterostructure.
+The sum of the excess charge on the PbTe and Pb side $\Delta Q$ is reported in the plot.
+Figure D7: BdG spectrum for the (4,-5) PbTe/Pb HS. The right panel displays the superconducting states dispersion in a narrower energy range.
+Figure D8: (a) SC–DOS $\rho(\varepsilon)$ and (b) ADOS $\chi(\varepsilon)$ for the (4,-5) PbTe/Pb HS computed with the fixed-$\Delta$ solution method.
+The triplet components of $\chi(\varepsilon)$ are clearly negligible.
+For the PbTe/Pb (4,-5) HS, we compared the components of $\rho(\varepsilon)$ and $\chi(\varepsilon)$ following the approach discussed for bulk Pb in Appendix 1.
+Using the fixed-$\Delta$ solution method, we observed a similar behaviour as for bulk Pb and concluded that the triplet components of the anomalous density can also be neglected in the HS (Fig.
+D8).
+Table D1: Excess of charge $\Delta Q$ computed for different strain.
 Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer.
@@
 9.0%/-0.4% -0.119 0.119
-Figure S11: Mulliken charge analysis for PbTe/Pb.
-Layer-by-layer excess charge.
-Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer.
-The red dashed line denote the separation between the PbTe and Pb side of the heterostructure.
-We notice that at the interface there is a depletion of charge on the PbTe side of the heterostructure, while the opposite is true on the Pb side.
-The sum of the excess charge on the PbTe and Pb side $\Delta Q$ is reported in the plot.